# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v2

@@ 2.5 Experimental datasets
$y=2x_{1}-x_{2}+x_{3}$ is used to generate a synthetic dataset which contains 3 features with values sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation $y=2x_{1}-x_{2}+x_{3}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this dataset to the language model, we fully specify the relationship between the features and targets to test the setting where the task description is fully descriptive.
@@
We then test four public benchmark datasets. The Breast Cancer dataset [36], where the task is to predict the diagnosis of a breast mass; the California Housing dataset [37], with the goal of predicting the median house value from US census data for a region; the Wine Quality dataset [38, 39] where the task is to predict the wine quality from its physicochemical characteristics; and the Heart Disease dataset [40, 41] where the task is to predict a heart disease diagnosis from physiological features.
@@ 3.1 Evaluating the LLM-elicited priors
Figure 2: Test accuracy of the posterior distribution for varied numbers of training data points. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. 2(a) shows the histogram of accuracy achieved whilst 2(b) shows a line plot of the average mean posterior accuracy or Mean Squared Error (MSE), with the error bars representing the 95% confidence interval, as well as the additional datasets tested in Appendix A.6. The green arrows point in the direction of improvement in the metric.
@@
For real-world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. On the other hand, for the Wine Quality dataset, Figure 2(b) shows that the prior distribution on $\theta$ elicited from the language model is too strict, as it eventually reduces the mean accuracy of the logistic regression model when the number of training points is larger than $25$, whilst the variation in accuracy remains smaller. Given these samples, the likelihood of the prior distributions could have been measured (as described in 2.4) before the elicited distribution was used for new predictions.
@@
Interestingly, for Heart Disease, there is no noticeable difference between the average accuracy of a logistic regression model under an uninformative or elicited prior. This is because the elicited prior contains many Gaussian distributions of mean 00 and standard deviation $1$, making it very similar to the uninformative prior. This can be seen in Figure 11 within Appendix A.7 which plots the elicited parameter distributions as a histogram in each feature dimension, for the first $10$ features. This is reassuring, since it shows that the language model is able to respond with an uninformative prior when it is not confident in its own prior knowledge of a task.
@@
For both the California Housing and UTI datasets, the elicited prior provides a greater mean accuracy over the uninformative prior for all of the number of observed data points tested, illustrating the potential of using a language model for prior elicitation. This should be of particular note considering that the UTI dataset is private and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that the use of prior knowledge from language models is most relevant when users have access to only a few samples.
@@
This can be important in scenarios where collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating model development and validation.
@@
In Appendix A.6, we present additional accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Wine Quality datasets, which supplement the results in Figure 2(b). We further test the number of task descriptions required in Appendix A.11, and the quality of those descriptions in A.10.
Given two sets of priors, one elicited from a language model and an uninformative prior, the prior predictive distribution can be sampled and the likelihoods calculated given some data points. This can be done in a way similar to that described in Section 3.6.
@@ 3.2 Incorporating expert information
However, it is interesting that without this expert information, the LLM had already correctly identified the relationship between the risk of UTI and the frequency of the bathroom, nighttime disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in posterior accuracy over LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
In Appendix A.10, we explore the use of including expert information in the task for a synthetic dataset, where we incrementally include more detailed information.
@@ 3.3 Extracting the in-context prior and posterior
This section explores the prior and posterior distributions of the linear predictive model that the LLM is using in-context. Further experimental details are available in Appendix A.5.3.
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (as outlined in Section 2.3.2) are shown in Figures 4 and 5. Once we estimate the parameters of the linear model, we can use them to calculate the logits (or regression values) of the predictions returned by the language model (from in-context learning) and measure the error between the outputs of the language model and the estimated internal model. This is shown in the bottom row of Figures 4 and 5, suggesting that our estimates of the previous internal parameters are reliable for the real-world datasets. The upper row presents the distributions of the estimated parameters, with each colour representing a single feature.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model uses some knowledge about the tasks when making predictions, even without any demonstration examples. However, we see that the language model often uses a prior centred around 00, suggesting that it is not particularly informative.
The extracted posterior distributions are shown in Figure 5 and show some differences compared to the prior distributions. For California Housing, Heart Disease, and Wine Quality, our estimation error increased but remained low. However, we find that the Breast Cancer dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the internal model parameters are not reliable for this case. It is interesting that, on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution, and this is likely due to the limitations of the language model and in particular, its capacity to understand demonstrations and instructions in the same prompt.
We do not estimate the language model’s posterior distribution for the UTI dataset as we keep the samples in this dataset private and are therefore unable to provide demonstrations in-context.
@@ 3.4 Is the language model being truthful?
UTI Breast Cancer California Housing Wine Quality Heart Disease
0.39 0.31 0.19 0.13 0.18
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model uses a different prior distribution to make its own predictions compared to the one it provides for us to construct a separate predictive model.
@@ 3.5 Is in-context learning approximating Bayesian inference?
Breast Cancer California Housing Wine Quality Heart Disease
0.01 ± 0.00 0.32 ± 0.09 0.02 ± 0.00 0.04 ± 0.01
It is important that the model performs Bayesian inference, so that newly observed data points and prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context, by measuring the difference between the extracted in-context posterior distribution and Monte Carlo sampling applied to the extracted prior in-context distribution (Section 2.3.4).
The results of this analysis are given in Table 2, which shows to what extent the posterior distribution extracted from the language model’s in-context model agrees with Monte Carlo sampling on the extracted in-context prior.
For California Housing, we find that we can reliably approximate the in-context prior and posterior models but find a large difference between the sampled posterior (using Monte Carlo methods) and the extracted posterior (using MLE sampling). For this task, the language model failed to perform Bayesian inference in-context, and may suggest the use of prior elicitation instead.
However, for Wine Quality and Heart Disease, the difference between the extracted posterior and the posterior samples of the extracted prior matches closely, and the error in our estimations of the parameters is low (Figures 4 and 5). In these cases, we are more confident that in-context learning is imitating Bayesian inference, and so should use a model selection tool to decide between prior elicitation and in-context learning.
Appendix A.8 presents histograms of the parameter values of each feature after Monte Carlo sampling or in-context learning have been performed. Figure 12 within Appendix A.8 supports the findings in this section and allows us to visualise the differences illustrated by Table 2.
These results provide empirical evidence that in-context learning sometimes approximates Bayesian inference, but not necessarily for all tasks.
@@ 3.6 Bayes factor and model selection
Figure 6: Distribution of the prior accuracy or mean squared error, the marginal likelihood, and the Bayes factor. The top row shows the distribution of prior accuracy, whilst the bottom row shows the distribution of prior likelihoods. These were calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is then calculated between prior elicitation and in-context learning, with a number larger than $1$ favouring prior elicitation. Note that for $y=2x_{1}-x_{2}+x_{3}$ and California Housing (regression tasks), results are plotted on the log-scale, and we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles.
To compare prior elicitation with in-context learning (when in-context learning is approximating Bayesian inference), we can calculate the Bayes factor as in Section 2.4.
In the case of prior elicitation, we test the model as described in Section 2.2 and Equation 1. We approximate the log-likelihood of this model by sampling the prior predictive log-likelihood of $\mathcal{D}$. For this, we use the same prior distribution $p_{M,T}(\theta)$ presented in 3.1 and sample $500$ predictions for $25$ data points, repeated with $5$ different splits of the dataset, to get samples of $p(\mathcal{D}|\alpha_{0})$
@@
Figure 6 shows the distribution of prior accuracy and marginal likelihood for in-context learning, prior elicitation, and an uninformative prior along with the calculated Bayes factor.
For the Breast Cancer and Wine Quality datasets, prior elicitation is preferred over in-context learning, whereas for the Heart Disease dataset, in-context learning performs better. In the case of Breast Cancer and Wine Quality, this is likely because the prior distributions elicited from the language model were more informative compared to the one that the language model used in-context. For Heart Disease, the opposite is true, since when performing prior elicitation, the language model returned an uninformative prior (Section 3.1), whilst it used a prior with a smaller parameter range in-context (Section 3.3).
For regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for California Housing, in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and low marginal likelihood, making the Bayes factor unreliable to compute. Furthermore, in the case of California Housing, in-context learning is not reliably estimating Bayesian inference (Section 3.5), suggesting that prior elicitation may still be preferable.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with both methods being the better choice over a uninformative prior for all cases tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
@@ 3.7 Choosing between priors
As seen in Appendices A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this might not be the case for all datasets or language models. Given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian information criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Furthermore, this method also allows us to compare the elicited distributions to the uninformative prior to ensure that the priors from our language models do not harm predictive performance. By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
@@ 4 Limitations and future research
In this work, we treat the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could be asked to produce more detailed task descriptions (Section 3.2 and Appendix A.10), but further tasks would be interesting to explore.
@@
An interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Further exploration of the methods would be beneficial.

### v3

@@ 2.5 Experimental datasets
The equation $y=2x_{1}-x_{2}+x_{3}$ is used to generate a synthetic dataset that contains 3 features sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation and perturbed by noise $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this dataset to the language model, we fully specify the relationship between the features and targets to test the setting where the task description is fully descriptive.
@@
We then test four public benchmark datasets. The Heart Disease dataset [36, 37] where the task is to predict a heart disease diagnosis from physiological features; the Diabetes dataset [38], with the goal of predicting a quantitative measure of disease progression one year after baseline; the Hypothyroid dataset [39, 40], where the task is to predict whether a patient has hypothyroidism based on biomarkers; and the Breast Cancer dataset [41], where the task is to predict the diagnosis of a breast mass.
@@ 3.1 Evaluating the LLM-elicited priors
Figure 2: Test accuracy of the posterior distribution for varied numbers of training data points. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For Diabetes, a regression task, we report the mean squared error on log-scale. 2(a) shows the histogram of accuracy achieved whilst 2(b) shows a line plot of the average mean posterior accuracy or Mean Squared Error (MSE), with the error bars representing the 95% confidence interval, as well as the additional datasets tested in Appendix A.6. The green arrows point in the direction of improvement in the metric.
@@
Figure 2(b) shows that for real-world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. The elicited prior distribution achieved a greatly increased accuracy for lower numbers of observed data points, enabling more reliable predictions earlier into model development.
In the case of Hypothyroid prediction, using the elicited priors did not improve the posterior predictions for low numbers of observed data points. However, once the number of observed data points increased to $20$ and greater, the accuracy was higher when using the elicited priors. As in Section 2.4 and 3.6, with a Bayesian approach, the likelihood of the prior distributions can be measured given the observed data before they are used for new predictions and the appropriate prior can be selected without access to a test set.
@@
Interestingly, for Heart Disease, there is no noticeable difference between the average accuracy of a logistic regression model under an uninformative or elicited prior. This is because the elicited prior contains many Gaussian distributions of mean 00 and standard deviation $1$, making it very similar to the uninformative prior. This is illustrated in Figure 12 within Appendix A.7 which plots the elicited parameter distributions as a histogram in each feature dimension, for the first $10$ features. This is reassuring, since it shows that the language model is able to respond with an uninformative prior when it is not confident in its own prior knowledge of a task.
@@
For both the Diabetes and UTI datasets, the elicited prior provides a greater mean accuracy or lower MSE over the uninformative prior for all of the number of observed data points tested. This should be of particular note considering that the UTI dataset is private and will have never appeared within the language model’s training.
@@
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that the use of prior knowledge from language models is most relevant when users have access to only a few samples. This can be important in scenarios where collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating model development and validation.
@@
In Appendix A.6, we present additional accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Hypothyroid datasets, which supplement the results in Figure 2(b). In Appendix A.11, we examine the number of task descriptions required to elicit consistent prior distributions, and how the detail of these descriptions affects posterior performance in A.10.
@@ 3.2 Incorporating expert information
It should be noted that without this expert information, the LLM had already correctly identified the relationship between the risk of UTI and bathroom usage, the frequency of nighttime disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in posterior accuracy over LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
In Appendix A.10, we examine the elicited priors when incrementally including more detailed expert information in the task description for a synthetic dataset, which a language model holds no prior knowledge for.
@@ 3.3 Extracting the in-context prior and posterior
This section explores the prior and posterior distributions of the linear predictive model that the LLM is using in-context (Appendix A.5.3).
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (Section 2.3.2) are shown in Figures 4 and 5. To quantify our approximation error, once we estimate the parameters of the linear model, we can use them to calculate the predictions returned by the language model in-context to measure the difference between the output of the language model and the estimated internal model. This is shown in the bottom row of Figures 4 and 5, suggesting that our estimates of the internal parameters are reliable for the real-world datasets.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model uses some knowledge about the tasks when making predictions, even without any demonstration examples. However, we see that the language model often uses a prior centred close to 00, suggesting that it is not particularly informative.
The extracted posterior distributions are presented in Figure 5 and show some differences compared to the prior distributions. For the Diabetes and Heart Disease tasks our estimation error increased but remained low. However, we find that the Breast Cancer and Hypothyroid dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the internal posterior model parameters are not reliable for this case. The error in approximating the parameters of the in-context model for the synthetic task is discussed in Appendix A.3. It is interesting that, on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution. This is likely due to the limitations of the language model, and in particular, its capacity to understand demonstrations and instructions in the same prompt.
As LLMs improve their ability to follow instructions, these approximation errors will reduce considerably since they are a product of the language model’s inconsistent imitation of a linear model.
We do not estimate the language model’s posterior distribution for the UTI dataset as samples in this dataset are kept private, and so we are unable to provide demonstrations in-context.
@@ 3.4 Is the language model being truthful?
UTI Heart Disease Hypothyroid Diabetes Breast Cancer
0.39 0.18 0.15 0.38 0.31
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model uses a different prior distribution to make its own predictions, to the one it provides for us to construct a separate predictive model.
@@ 3.5 Is in-context learning approximating Bayesian inference?
It is important that the model performs Bayesian inference, so that newly observed data points and prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context. This is done by measuring the difference between the extracted in-context posterior distribution and Monte Carlo (MC) posterior samples of the extracted prior in-context distribution (Section 2.3.4).
Figure 6: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset we have removed the long tails for improved visualisation.
The results of this analysis are given in Figure 6, which shows the extent to which the posterior distribution extracted from the language model’s in-context model agrees with Monte Carlo sampling on the extracted in-context prior. In this figure, we present the results for those datasets which achieve a low approximation error when estimating the posterior in-context parameters (Figure 5). The results of the other datasets are given in Appendix A.8, alongside those on other dataset splits.
The results for the two datasets given in Figure 6 present contradictory conclusions. For the Diabetes task, the posterior parameter samples from the in-context posterior and MC sampling on the in-context prior visually agree and produce a low energy statistic, suggesting that the LLM was able to approximate Bayesian inference in-context. On the other hand, for the Heart Disease task, we find that the two methods of sampling posterior parameters yielded different results, implying that for this task, the LLM was not correctly approximating Bayesian inference.
These results provide empirical evidence that whether in-context learning approximates Bayesian inference depends on the task. The LLM inconsistently applies the Bayes rule in context, which suggests that it should not be considered a reliable classifier. In this case, our prior elicitation framework presents itself as the more dependable option.
As the field develops, it is feasible that future LLMs can reliably perform Bayesian inference, and so in these cases we require a model selection tool to decide between different methods for incorporating a language model’s knowledge into a predictive task.
@@ 3.6 Bayes factor and model selection
Figure 7: The distribution of the prior marginal likelihood, and the Bayes factor. This shows the distribution of prior likelihoods for the tasks. These were calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is then calculated between prior elicitation and in-context learning, with a number larger than $1$ favouring prior elicitation. For the regression tasks, we use the log-scale and exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles.
To compare prior elicitation with in-context learning, when in-context learning is approximating Bayesian inference, we can calculate the Bayes factor as described in Section 2.4.
In the case of prior elicitation, we test the model as described in Section 2.2 and Equation 1. We approximate the marginal log-likelihood of this model by sampling the prior predictive log-likelihood of $\mathcal{D}$. For this, we use the same prior distribution $p_{M,T}(\theta)$ presented in 3.1 and sample $500$ predictions for $25$ data points, repeated with $5$ different splits of the dataset, to get samples of $p(\mathcal{D}|\alpha_{0})$
@@
Figure 7 shows the distribution of the marginal likelihood for in-context learning, prior elicitation, and an uninformative prior along with the calculated Bayes factor (where reliably defined).
In the case of predicting Breast Cancer and Hypothyroidism, prior elicitation achieves a greater mean likelihood than both an uninformative prior and in-context learning. For these tasks, the Bayes factor and distribution of likelihoods suggest the use of prior elicitation, since it better describes the data (in terms of likelihood) and is cheaper to use for inference.
The Heart Disease prediction task favoured in-context learning, which provided a higher mean likelihood. For this task, the LLM returned an uninformative prior for the elicited distribution (Section 3.1), and used a prior with a smaller parameter range in-context (Section 3.3). In this case, the final choice of approach should depend on the setting, as using in-context learning for predictions might be too costly or unrealistic. Unlike the Bayes factor, the Bayesian Information Criterion (BIC) technique accounts for the computational complexity of classifiers (Section 2.4), but due to the large parameter sizes of an LLM and the availability of APIs, it could unfairly disfavour in-context learning. This approach may be more informative for some settings, and for Heart Disease prediction, would recommend prior elicitation.
For regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for Diabetes, in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and so a low marginal likelihood, making the Bayes factor unreliable to compute. Again, in the case of Diabetes prediction, the final model decision relies on the available resources.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with prior elicitation being the better choice over a uninformative prior for all datasets tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
@@ 3.7 Choosing between priors
As seen in Appendices A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian Information Criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Furthermore, this method also allows us to compare the elicited distributions to the uninformative prior to ensure that the priors from our language models do not harm predictive performance. By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
@@ 4 Limitations and future research
In this work, we consider the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could work alongside an LLM by producing more detailed task descriptions (Section 3.2 and Appendix A.10), but other ways of integrating experts would be interesting to explore.
@@
Finally, an interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Another possible solution is to use the output of in-context predictions as additional features to be learnt from. Further exploration of methods in this direction would be beneficial.